RC Meeting, Tuesday, January 18, 2022

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 AM, remote

  • Zoom remote connection information:
 Join ZoomGov Meeting
 Meeting ID: 160 710 0885
 Passcode: 595071
 One tap mobile
 +16692545252,,1607100885# US (San Jose)
 +16468287666,,1607100885# US (New York)
 Dial by your location
       +1 669 254 5252 US (San Jose)
       +1 646 828 7666 US (New York)
       +1 669 216 1590 US (San Jose)
       +1 551 285 1373 US
       833 568 8864 US Toll-free
 Meeting ID: 160 710 0885
 Find your local number: https://jlab-org.zoomgov.com/u/ab2CekLXgQ

Useful Links


  • Social distancing requirements back in place. Please maintain a separation of 6 feet during shifts.
  • See updated guidance on JLab COVID-19 isolation and quarantine protocols: https://www.jlab.org/memo/jlab-covid-19-isolation-and-quarantine-requirements
    • Communicate with Occupational Medicine if you have COVID-19 symptoms, have tested positive for COVID-19, or have been exposed to someone who has tested positive for COVID-19: (jnewman@jlab.org or jbanks@jlab.org).
  • Please keep an eye on the HMS NMR during shift. Due to the NMR locking instability, the alarm has been been disabled, so slow drifts of the NMR signal may be missed.
  • See note on lower right TV screen for bringing up Zoom in CH for remote 3rd shifts

Current Status

Current RC: Dave Gaskell

Future plans

  • Complete LH2 (center)
  • Rotate SHMS to 8.39 degrees for SHMS left setting

Current Run Plan

  • Complete Q2=3.85, W=3.07, low epsilon, SHMS center data taking on LH2.
    • SHMS settings: theta = 6.385 degrees, P=6.538 GeV
    • HMS settings: theta = 35.12 degree, P=-1.332 GeV
    • Targets = LH2
  • Change to SHMS left --> theta = 8.39 degrees
    • Targets = LH2,aluminum dummy, LD2

Notices in Effect for Shift Crew

  • The Raster Coil pair Ax and Ay are off. Raster needs to be at 4x4. Raster gui has been updated - can now request 2x2.
  • With LHe alarms it is good to check the cameras for plumes. If no plume and it is below the LO alarm and above LOLO then monitor with a plot and halog the problem.
  • With LNitrogen alarms are not that critical. Just need to start monitoring the problem with plot and Halog.
  • HMS NMR is not locking stably - the alarm is turned off. Please check on the NMR readback (while it is locked) intermittently to make sure the HMS dipole field is not drifting,

Current Issues

  • Vertical beam position on target has changed a little: Dave Meekins is not concerned about an offset of this size (<0.5 mm).
  • There is a rotation issue that has come up if moving HMS between 19 to 25 degrees. It usually gets stuck around this range, please be wary of this issue. We will likely skip settings in this range if this drive controller error cannot be solved. For singles it will not be a big deal to skip settings for HMS at those angles. We need a longer access to address this issue.
  • The Raster Coil pair Ax and Ay are off
  • Diffuser limit at 46uA (Note: is this still valid?)

Crew Chiefs, We have permission to raise the Hall C diffuser raster limit to allow running >46uA to Hall C. This limit is normally imposed when one of the two rasters is not operative. Since the hall was able to double the running raster current to get the same size raster on the target as having both rasters, Keith is allowing us to raise the limit. To get to the setting, open the Hall C diffuser interlocks screen. Go to the bottom right corner and open the Expert screen. Just below and to the right of the FSD bits, raise the "Raster 1 OK Limit" to slightly above the highest current the hall wants to run at. Do not change the Raster 2 limit. Make a log entry with screen shots. Make an ABIL entry, System MPS and an expected removal date of 3 Jan. 2022. Paul

Planned Accesses to the Hall

  • None

Planned Accelerator Configuration Changes

  • Hall B pass change planned for Monday, Jan. 24

Special Requests

  • None